### v1

# Evaluating Mathematical Reasoning Beyond Accuracy

Abstract
The leaderboard of Large Language Models (LLMs) in mathematical tasks has been continuously updated. However, the majority of evaluations focus solely on the final results, neglecting the quality of the intermediate steps. This oversight can mask underlying problems, such as logical errors or unnecessary steps in the reasoning process. To measure reasoning beyond final-answer accuracy, we introduce ReasonEval, a new methodology for evaluating the quality of reasoning steps. ReasonEval employs validity and redundancy to characterize the reasoning quality, as well as accompanying LLMs to assess them automatically. Instantiated by base models that possess strong mathematical knowledge and trained with high-quality labeled data, ReasonEval achieves state-of-the-art performance on human-labeled datasets and can accurately detect different types of errors generated by perturbation. When applied to evaluate LLMs specialized in math, we find that an increase in final-answer accuracy does not necessarily guarantee an improvement in the overall quality of the reasoning steps for challenging mathematical problems. Additionally, we observe that ReasonEval can play a significant role in data selection. We release the best-performing model, meta-evaluation script, and all evaluation results at https://github.com/GAIR-NLP/ReasonEval.
1 Introduction
Mathematical reasoning, a core cognitive skill, is crucial for resolving complex problems and making informed decisions (Hendrycks et al., 2021; Lewkowycz et al., 2022), playing a significant role in large language models (LLMs) research (Azerbayev et al., 2023; Lu et al., 2023). Given its significance, reliably evaluating mathematical reasoning in LLMs becomes crucial. Current methodologies to evaluate mathematical reasoning in LLMs focus primarily on the final result (Luo et al., 2023; Chern et al., 2023; Yu et al., 2023), neglecting the intricacies of the reasoning process. For example, the OpenLLM leaderboard,111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard a relatively well-recognized benchmark for LLMs, uses overall accuracy to assess models’ mathematical reasoning. Despite being effective to some degree, such evaluation practice could mask underlying issues such as logical errors or unnecessary steps that compromise accuracy and efficiency of reasoning steps. In this work, we argue that a desirable evaluation criterion for mathematical reasoning encompasses not only the accuracy of the final answer but also the correctness and efficiency of each step in the reasoning process. Moreover, it is imperative that the evaluation metrics be open-source and replicable to ensure transparency and reliability.
Our design philosophy stems from the fact that a correct final answer does not guarantee a flawless reasoning process (Lewkowycz et al., 2022; Uesato et al., 2022), and excessive or irrelevant reasoning steps can lead to potential errors as well as increased computational costs (Zhang et al., 2023). Once these issues go unnoticed, they can cause problems in many application scenarios. For example, in K12 mathematics education, incorrect or redundant solution steps provided by LLMs could mislead students. There have been some recent works related to the above evaluation principle. Specifically, Uesato et al. (2022); Lightman et al. (2023); Wang et al. (2023) present process reward models for mathematical reasoning,222 Uesato et al. (2022) and Lightman et al. (2023) are close-source. which focus on their utility as verifiers (Cobbe et al., 2021) to boost the accuracy (i.e., by generating many candidate solutions and selecting the one ranked highest by the verifier), leaving its underlying potential to identify reasoning errors333Identifying reasoning errors and ranking reasoning steps are different tasks. As also highlighted by Tyen et al. (2023), even state-of-the-art (SOTA) models like GPT-4 (OpenAI, 2023) still struggle to detect errors in reasoning tasks. and redundancy. Clinciu et al. (2021); Golovneva et al. (2022) rely on embedding-based methods to measure the quality of general reasoning explanation, which has limitations in handling diverse reasoning steps from various models.
In response to these challenges, we propose ReasonEval, a suite comprising a new evaluation methodology with defined metrics for assessing mathematical reasoning quality and corresponding LLM-based evaluators for automated calculation. As illustrated in Figure 1, ReasonEval emphasizes the validity (i.e., the step contains no mistakes in calculation and logic) of each reasoning step, and evaluates the redundancy (i.e., the step lacks utility in solving the problem but is still valid) of steps to ensure efficiency. We explore different design options for the LLM-based evaluators and show training such a high-quality evaluator for mathematical reasoning is not a straightforward task, which calls for an LLM with strong mathematical knowledge as well as specialized datasets annotated in accordance with our evaluation criteria. With these prepared, ReasonEval achieves SOTA performance on human-labeled datasets (§4.1) and can accurately detect different types of errors generated by perturbation (§4.2).
Furthermore, we show the utility of ReasonEval through the lens of two preliminary applications, as detailed in §5: (1) evaluating different LLMs trained for mathematical reasoning; and (2) selecting high-quality data to train such LLMs. We have the following main findings: (1) We find that an improvement in the result accuracy is not sufficient to ensure an enhancement in the overall quality of reasoning steps in challenging mathematical problems; (2) We find that the model scale, the base model, and the training methods have significantly influenced the quality of reasoning steps; (3) We find that ReasonEval can select high-quality training data to improve the efficiency of solving problems and the quality of solutions. These findings pave the way for future work to design methods that take into account the process of problem-solving.
Overall, our main contributions are as follows:
(1) We recognize the potential gap between what we are evaluating and what we are desiring in mathematical reasoning, prioritize validity and redundancy aspect to address the misalignments and inefficiencies that can arise in mathematical reasoning processes (§3). (2) We design a meta-evaluation testbed to assess the reliability (§4) and utility (§5) of mentioned metrics, guiding us to a superior metric design, solidifying the foundation for future exploration and significantly lowering trial and error costs. (3) We open-source our best-performing metric, meta-evaluation script and all evaluation results to facilitate future research.
2 Preliminaries
2.1 Problem Formulation
Given a mathematical problem , solution steps and an answer generated by LLMs, the goal is to evaluate how well the generated solution and answer are. Usually, the ground truth answer is available but the reference solution step is not always provided.
2.2 Existing Evaluation Methodology
Answer-based Matching
Reference-based Scoring
Instead of only using the final result as a scoring criterion, some works (Sawada et al., 2023) try to measure the reasoning quality by comparing the similarity between generated and reference solution steps (i.e., and ). Although datasets like MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) provide the ground truth solutions, the existence of diverse reasoning paths leading to the same answer renders reliance on any single one of them as a reference unreliable.
Prompting-based Method
This method directly asks LLMs with a well-defined prompt to generate a judgment for a given generated solution and answer (Dubois et al., 2023). This approach usually requires a very powerful LLM, often GPT-4, which may not be practical for iterative model development due to cost and transparency concerns.
The aforementioned methods either focus solely on the final results (e.g., ) or are limited by the use of reference solutions and proprietary LLMs, and most of them only concentrate on evaluating “correctness”, neglecting other aspects such as the redundancy of solution steps. This has inspired us to propose a set of evaluation criteria that better aligns with the reasoning models in the era of LLMs.
3 ReasonEval: Metrics and Implementations
3.1 Design Principle
We argue that for tasks involving multi-step reasoning, measuring the quality of reasoning should not solely depend on the correctness of the final result. It should also consider (1) the accuracy of each reasoning step; (2) the efficiency of the reasoning process. To this end, we design evaluators that assess reasoning steps regarding validity and redundancy in a step-by-step format, checking whether each step advances well towards solving the problem. Precisely, validity denotes the step contains no mistakes in calculation and logic, and redundancy describes the step lacks utility in solving the problem but is still valid.
3.1.1 Task Formulation
Following this principle, we formulate such a metric design process as a three-way classification task. Given a question and the solution steps , a label { positive, neutral, negative } will be assigned with the following probability:
where the positive label indicates the step is correct and contributes to solving the question, while the neutral label represents that the step is also correct but does not make any progress. The negative label signifies an incorrect step.
3.1.2 Scoring scheme
We can define step-level scores for each step:
Assuming we have a set of scores for a solution containing N steps, We can define solution-level scores:
3.2 Evaluator Design
3.2.1 Model Architecture
LLM Backbone
Our defined evaluation method (Eq.1) can be implemented by directly prompting existing LLMs or fine-tuning LLMs using supervised dataset. To make a comprehensive study, in this work we instantiate ReasonEval by different types of LLMs, which vary in base model types (e.g., Llama2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023)), model sizes (e.g., 7B, 34B), and post-training strategies (e.g., continued pretraining (Azerbayev et al., 2023) and fine-tuning).
Task-specific Layer
Our evaluator’s architecture is identical to the base model, except that the classification head for next-token prediction is replaced with a classification head for outputting the possibilities of each class. After normalization with the softmax layer, we get the possibility for each class.
3.2.2 Fine-tuning Data
The above task formulation allows us to utilize the existing dataset, PRM800K (Lightman et al., 2023), as training data in ReasonEval. PRM800K contains about 800,000 step-level labels over 75,000 solutions. It is collected by employing humans to label the step-by-step solutions generated by GPT-4 to MATH problems. The labels for each step are identical to those mentioned in §3.1.1.
4 Experiment
4.1 Meta-Evaluations on Human Judgement Datasets
| Invalid | Redundant | |||||||
| Solution-level | Step-level | Solution-level | Step-level | |||||
| F1 Score | AUC | F1 Score | AUC | F1 Score | AUC | F1 Score | AUC | |
| Embedding-based Methods | ||||||||
| ROSCOE-SA | 48.2 | 57.5 | - | - | 50.7 | 53.9 | - | - |
| ROSCOE-SS | 51.6 | 49.6 | - | - | 52.0 | 52.7 | - | - |
| Prompting-based Methods | ||||||||
| GPT-3.5-turbo | 59.7 | - | 53.2 | - | 53.0 | - | 51.5 | - |
| GPT-4 | 73.2 | - | 61.0 | - | 57.1 | - | 54.2 | - |
| Step-level Evaluators | ||||||||
| Math-shepherd-Mistral-7b | 70.1 | 77.3 | 60.0 | 77.2 | 50.4 | 54.5 | 42.7 | 53.0 |
[…]
| ReasonEval | 72.8 | 81.9 | 67.7 | 83.9 | 60.5 | 65.6 | 59.0 | 68.3 |
| ReasonEval | 78.0 | 85.1 | 68.6 | 85.7 | 60.7 | 63.4 | 59.7 | 70.9 |
| ReasonEval | 74.7 | 84.3 | 76.6 | 90.5 | 59.6 | 63.0 | 58.6 | 68.3 |
| ReasonEval | 77.3 | 86.2 | 70.4 | 90.5 | 58.6 | 63.6 | 59.5 | 71.8 |
| ReasonEval | 78.6 | 87.5 | 73.9 | 89.5 | 61.6 | 64.8 | 59.7 | 72.2 |
| ReasonEval | 79.6 | 90.8 | 77.5 | 92.8 | 58.3 | 62.7 | 57.5 | 67.3 |
4.1.1 Meta-Evaluation Setup
Meta-evaluation Datasets
Inspired by Zeng et al. (2024), which propose a benchmark named Meta-Reasoning-GSM8K (MR-GSM8K), we construct a meta-evaluation dataset MR-MATH to better assess how well different evaluators can detect errors in reasoning steps. It is constructed as follows: (1) To collect the first type of errors affecting the correctness of steps, we recruit undergraduates who have a solid mathematical background to label solutions generated by Abel (Chern et al., 2023) and WizardMath (Luo et al., 2023). We collect 83 samples with incorrect steps and 76 without, pinpointing the first error location in the former. All the solutions reach the correct final answers, meaning the evaluators need to judge correctness based on the process rather than the outcome. Additionally, since the style of solutions differs from the training set of PRM800K (See examples of solutions in Appendix A.1), it tests the generalization of ReasonEval to some degree. (2) For the second type of errors affecting the efficiency of problem solving process, as they are more rarer than the first one, we sample solutions from the test set of PRM800K directly, containing 150 samples with redundant steps and 150 samples without.
Evaluators
We compare three methods to evaluate reasoning steps automatically: embedding-based methods, prompting-based methods and step-level evaluators. For the embedding-based methods, we choose ROSCOE (Golovneva et al., 2022), a SOTA method among them. Specifically, we select the semantic alignment metrics (ROSCOE-SA) and the semantic similarity metrics (ROSCOE-SS), which do not require references solution steps. For the prompting-based methods, we select GPT-3.5-turbo and GPT-4, two mainstream generation models. We use the prompt suggested by Zeng et al. (2024) for the invalid errors and the modified version for the redundant errors. For the step-level evaluators, we select 7 base models for ReasonEval: Llama-2-7B, Mistral-7B, Llemma-7B, Llemma-34B, WizardMath-7B-V1.0 (Luo et al., 2023), WizardMath-7B-V1.1, Abel-7B-002 (Chern et al., 2023). The llemma series is continuing pertaining on math-based corpus. The Abel and WizardMath series are fine-tuning for solving mathematical problems. We also select the open-source SOTA process reward model Math-shepherd-Mistral-7B (Wang et al., 2023) to compare. The settings for these methods are in Appendix A.3 and the training details for ReasonEval are in Appendix A.2.
Meta-evaluation Metrics
We test two abilities: judging whether the solution contains errors and locating the error step. In the first task the ground truth is labels for solutions (solution-level), and in the second tasks the ground truth is labels for steps (step-level). Since both are classification tasks, we present the macro F1 score as a performance metric. Additionally, for any methods that require setting a threshold, we report the Area Under the Curve (AUC) metric, which is a more balanced evaluation of performance across different threshold settings.
4.1.2 Results and Analysis
Overall Results
We present our main results in Table 1. ReasonEval can outperform all other methods across all error types and label levels. It is noteworthy that the identification of errors at the step level is generally more challenging than at the solution level for all methods. This suggests a higher complexity in pinpointing errors at the individual step level. Furthermore, identifying redundant errors is more intricate than invalid errors due to the inherent ambiguity involved in the former.
Figure 2: Correlation between Pass@1 for the base model on MATH and AUC for the solution-level labels.
Analysis on Base Models
We investigate the correlation between the base LLMs’ mathematical reasoning capabilities and the performance of ReasonEval fine-tuned from them. In Figure 2, for the invalid errors, an increase in AUC shows a positive correlation with the base models’ ability to solve MATH problems, indicating that enhancing mathematical problem-solving abilities is beneficial for identifying invalid errors. It is noted that the Llemma-34B outperforms all 7B models, although its Pass@1 is not the highest. It highlights the importance of model scale and the continued pretraining over math-related corpus. For the redundant errors, the distinction across different base models is small, and it does not show the correlation as the invalid errors. This may be due to developers for these models prioritizing the correctness of solutions over their efficiency, resulting in similar performances in this aspect across various LLMs.
| Dataset | # Solutions | Labels | Problem Source | Labeled Method |
| Math-shepherd | correct, incorrect | MATH, GSM8K | unsupervised | |
| PRM800K | positive, neutral, negative | MATH | human |
Analysis on Training Data
In Table 1, the Mistral-7B trained on Math-Shepherd falls behind that trained on PRM800K. We summarize the information about Math-shepherd and PRM800K in Table 3. For Math-shepherd, although having 6x training data, there is more noise in the step-level labels generated with the unsupervised method, thus harming its precision in identifying errors. And the only two classes of labels also make it limited in evaluating redundancy.
Generalization Analysis
To assess the out-of-distribution generalization of ReasonEval, we evaluate its performance on MR-GSM8K, which consists of problems from GSM8K. We present the results in Table 2. We observe that both and achieve superior performance at the step level and approach the performance of GPT-4 at the solution level, demonstrating strong generalization capabilities. For Math-shepherd-Mistral-7B, it performs best in solution-level labels but lags behind in step-level labels, suggesting that the noise in step labels from Math-Shepherd negatively impacts its ability to accurately locate error steps.
4.2 Meta-Evaluations by Perturbation
To further analyze the effect of different types of errors on ReasonEval, we follow Golovneva et al. (2022) and introduce errors in the correct reasoning steps.
Setup
We sample 184 solutions without any invalid or redundant errors from the test set of PRM800K. These solutions cover various difficulty levels and subjects. We introduce 6 types of errors to one of the intermediate steps. In Table 4 we summarize the formats of the errors. All solutions are scored by ReasonEval and Math-shepherd-Mistral-7B.
Figure 3: Box-and-whisker plots of scores when meeting different types of errors at the step level. The boundaries of the whiskers are based on the 1.5 interquartile range.
Results
Figure 3 presents the results. We observe the scores for the first error step (step-level scores). We observe that: (1) The validity score exhibits a significant decrease while the redundancy score remains unchanged when meeting the logical and calculation errors (random number, random operation, swap steps, remove step). (2) For the repeated step, the validity score remains high and the redundancy score increases significantly, consistent with the fact that the repetition introduces redundancy to the overall solution without impacting its validity. (3) In the case of swapping solutions, where the solution is correct but not for the given problem, the validity score decreases and the redundancy score increases. It indicates ReasonEval treats the problem and the solution as a whole, rather than assessing the solution in isolation. (4) For Math-shepherd-Mistral-7B, it shows insensitivity to the errors causing redundancy like repetition step. In light of these observations, we conclude that ReasonEval can detect different types of errors, effectively distinguishing between those that affect the solution’s validity and those that introduce redundancy. We present other analysis and the results of solution-level scores in Appendix A.4.
5 Utilizing ReasonEval for Evaluating and Improving Reasoning Quality
5.1 Evaluating Reasoning Quality of LLMs Specialized in Math
Setup
The models selected for measurement are two mainstream LLMs specialized in math: Abel (Chern et al., 2023) and WizardMath (Luo et al., 2023), with different scales. We also report the results of LLaMA-2 (Touvron et al., 2023) naive fine-tuned on PRM800K (approximately 6K solutions that reach the correct final answers) for comparison. We evaluate the performance of the LLMs on MATH and sample 100 solutions for each problem to reduce evaluation noise and sampling randomness. Specifically, the sampling temperature is set to 0.6, and the top-p value is set to 0.95. The maximum sample length is set to 2048 tokens. All solutions are scored by ReasonEval.
Our analysis includes two aspects: (1) False positive rate: it refers to the proportion of solutions that have the correct final answers but contain incorrect steps among all solutions with correct final answers. (2) The redundancy of solutions. We also analyze the first error location in reasoning steps and present the results and analysis in Appendix B.1.
5.1.1 False Positive Rate
| Model | Base Model | Methods | Acc. (%) | FPR (%) |
| LLaMA2-13B-PRM800K | LLaMA2 | SFT | 7.4 | 40.6 |
| LLaMA2-70B-PRM800K | 14.7 | 21.8 | ||
| Abel7B-001 | LLaMA2 | SFT | 12.4 | 31.8 |
[…]
| Abel70B | 25.7 | 20.0 | ||
| Abel7B-002 | Mistral | 30.0 | 22.8 | |
| WizardMath7B-V1.0 | LLaMA2 | SFT+RLHF | 8.9 | 34.4 |
[…]
| WizardMath70B | 18.7 | 16.5 | ||
| WizardMath7B-V1.1 | Mistral | 31.0 | 16.7 |
Figure 4: Correlation between the final-answer accuracy and the false positive rate.
Figure 5: Box-and-whisker plots of (upper) and (lower). The boundaries of the whiskers are based on the 1.5 interquartile range.
First, we compare the validity scores of reasoning steps from different models in Figure 5. For solutions with the correct answers but wrong steps, the validity scores should behave similarly to solutions with the incorrect answers. So we can use it to estimate the threshold and calculate the false positive rate. A rough estimation for the threshold is between 0.15 and 0.35. To further ensure a relatively accurate estimation, we check the false positive rate of Abel13B and WizardMath13B by humans, sampling one solution for each problem. Combining above information, the threshold is set to 0.25. We present the main results in Table 5 and Figure 4.
Increased final-answer accuracy does not inherently ensure a corresponding reduction in the false positive rate. In Figure 4, it is clear that the false positive rate drops and stays at a level of about 20% as the final-answer accuracy rises. When comparing WizardMath7B-V1.1 with WizardMath70B, although the former exhibits a much higher accuracy (31.0% vs. 18.7%), its false positive rate is almost the same (16.7% vs. 16.5%). It indicates there exists a bottleneck for the quality of reasoning steps to advance.
The model size and base model affect the false positive rate significantly. When comparing LLaMA13B-PRM800K with LLaMA70B-PRM800K, although using the same training data, the distinction in false positive rates is large (40.6% vs. 21.8%). It indicates the importance of the model scale. The base model of Mistral (Jiang et al., 2023) also achieves a lower false positive rate than LLaMA2 (22.8% vs. 31.8%, 16.7% vs. 34.4%). Overall, the scaling law still applies to this field.
RLHF helps lower the false positive rate when the LLMs are relatively strong. There’s no distinction in false positive rates between SFT and SFT plus RLHF when the model is relatively weak, like 7B and the 13B model of LLaMA-2 (31.8% vs. 34.4%, 29.8% vs. 31.3%). As the model size becomes larger, SFT plus RLHF performs better (16.5% vs. 20.0%, 16.7% vs. 22.8%). This indicates that effective supervision by RLHF requires a strong mathematical ability of the model itself.
5.1.2 Redundancy of Solutions
We analyze the redundancy in reasoning steps. Our main results are in Figure 5. It is noteworthy that the redundancy scores of Abel family are usually higher. This is because the solutions from Abel often involve restating the problem first, which is considered a meaningless action towards solving the problem by our evaluator. Moreover, the redundancy scores for solutions that fail to reach the correct answers are higher. This suggests that when a model is unsure about how to solve a problem, it tends to make more attempts that lack meaningful progression.
5.2 Improving Reasoning Quality by Selecting Training Data
In this part, we explore the potential of ReasonEval to select high-quality training data for SFT.
| Filter | #Training | Acc. | Val. | Red. | #Token |
| - | 100% | 22.2 | 65.2 | 27.4 | 723.4 |
| val. | 76.7% | 22.0 | 65.9 | 26.4 | 699.9 |
| random | 76.7% | 20.1 | 62.5 | 27.4 | 765.6 |
| red. | 71.9% | 21.8 | 65.6 | 22.1 | 681.5 |
| random | 71.9% | 20.3 | 62.3 | 28.0 | 746.1 |
| red. & val. | 56.7% | 22.0 | 67.8 | 22.5 | 701.2 |
| random | 56.7% | 20.0 | 62.1 | 27.6 | 739.5 |
Setup
The MMIQC (Liu & Yao, 2024) dataset consists of a mixture of processed web data and synthetic question-response pairs used to enhance the reasoning capabilities of LLMs. We randomly sample 10K unique responses generated by GPT-3.5-turbo on MATH from the dataset. We then filter it using , specifically removing samples with validity scores below 0.5 or redundancy scores above 0.15. We also combine these two conditions. We SFT the mistral-7b in different subsets with 3 epochs and observe the performance on MATH. To compare, we randomly sample the same amount of training data three times and report the average performance.
Results
We make the following observations from Table 6: (1) In terms of accuracy, all groups filtered by outperform the random group and achieve performance close to that of the full dataset. (2) The average number of tokens for the solutions decreases in groups filtered by , indicating its advantage in increasing problem-solving efficiency. (3) The group that combines the two filtering conditions significantly improves the quality of reasoning steps with only about half the training data.
6 Related Work
Evaluating reasoning steps automatically
The technique used to assess reasoning steps automatically can be broadly divided into three groups: (1) Embedding-based methods: Golovneva et al. (2022) propose several metrics and uses the embedding-based calculation among hypothesis steps, reference steps, and problems to represent them; (2) Parsing-based methods: this approach aims to parse steps into structured forms, like ‘subject-verb-object’ frames (Prasad et al., 2023) or symbolic proofs (Saparov & He, 2022). However, it presents challenges for complex datasets such as MATH due to the intricate logic involved; (3) Prompting-based methods: due to the generalization of LLMs, given tailored prompts to SOTA LLMs, they can check the solutions and find the potential errors (Tyen et al., 2023; Zeng et al., 2024).
Process reward model (PRM)
Process reward models are mainly used in reinforcement learning to give feedback to LLMs to align with human logic in mathematics. Lightman et al. (2023) and Uesato et al. (2022) both evaluate the performance of PRM as a verifier against the outcome reward model. Ma et al. (2023) combine a check-and-generation idea with PRM to generate a more accurate response. Wang et al. (2023) propose a way to automatically construct process-wise supervision data. These studies focus on leveraging PRM to enhance accuracy. In contrast, our research explores the potential of PRM to develop new evaluation methods in mathematics, moving beyond mere accuracy improvements.
7 Conclusion
In this work, we develop ReasonEval, a new metric to evaluate the quality of reasoning steps from the perspective of correctness and efficiency. Our experiments show that ReasonEval can effectively detect different types of errors and achieves competitive performance compared to other methods. With ReasonEval, we find the inconsistency between final-answer accuracy and the quality of reasoning steps. Additionally, we discover its efficacy in selecting training data.
Appendix A Supplements for Meta-Evaluations
A.1 Examples of Solutions Generated by Different LLMs
We present examples of solutions generated by different LLMs in Figure 9, including LLaMA-PRM800K, Abel and WizardMath.
A.2 Training Details for ReasonEval
The ReasonEval model is trained on NVIDIA A100 GPUs (4 GPUs for the 7B models, 16 GPUs for the 34B models) using the supervised fine-tuning (SFT) framework provided by Dai et al. (2023). Specifically, the model is trained in 1 epoch. We set the batch size to 64 and maximum sequence length to 2048. We use a peak learning rate 1e-6 with 8 warmup steps and cosine learning rate decay to 0. We use AdamW (Loshchilov & Hutter, 2018) as our optimizer with and weight decay of 0.1. Only the last token of a step is taken into account in the loss function, as it contains the full information about the step.
A.3 Settings for Evaluators
The settings for evaluators are as follows: (1) Embedding-based methods: for ROSCOE we use the embedding layer “roscoe-512-roberta-base”. We select Repetition-token matrix for ROSCOE-SA and the Repetition-step matrix for ROSCOE-SS. The threshold is set to 0.025. (2) Prompting-based methods: we present the prompts in Figure 6. (3) Step-level evaluators: for the validity scores, the threshold is set to 0.5, and for redundancy scores, it is set to 0.15. Both are applied to the solution-level labels and step-level labels. For math-shepherd we set the threshold to 0.5 for all errors.
| The Prompt for identifying invalid errors | The Prompt for identifying redundant errors |
| Question: {question} Student Solution: {solution} Your task involves three parts: 1. **Step-by-step Evaluation:** Go through the student solution carefully and identify key errors and potential misunderstandings that led to the incorrect solution. 2. **Final Judgement:** Provide an overall judgement on the correctness of the student’s solution. 3. **First Error Step:** If the solution is incorrect, generate the step number where the first error occurs, otherwise generate N/A here Here’s the format I want: Step-by-step Evaluation: [Provide a step by step examination of the student solution and identify key errors and misunderstandings here.] Final Judgement: [Insert only **correct** or **wrong** here] First Error Step: [Insert either N/A or the step number where the first error occurs] Please follow this format without any additional introductory or concluding statements. | Question: {question} Student Solution: {solution} Your task involves three parts: 1. **Step-by-step Evaluation:** Carefully review the student’s solution. Identify any neutral steps that, while reasonable, do not offer new insight, advance the solution, or suggest a next step. 2. **Final Judgment:** Indicate whether the solution contains or does not contain any neutral steps. 3. **Neutral Steps Identified**: If neutral steps are present, list their numbers in list format; otherwise, insert N/A. Here’s the format I want: Step-by-step Evaluation: [Provide a step-by-step examination of the student solution and identify the neutral steps.] Final Judgment: [Insert only the word **Present** if neutral steps are identified, or **Absent** if not.] Neutral Steps Identified: [Insert either N/A or the step numbers in list format.] Please follow this format without any additional introductory or concluding statements. |
A.4 Addition Information for Meta-Evaluations by Perturbation
In Figure 3, for swapping solutions, the redundancy score of the original solution is higher than others. This observation may stem from the fact that the first step (i.e., the observed step) typically involves analyzing the problem from scratch, making it more likely to involve restating the problems, which may be considered an unproductive or unnecessary step for the overall solution.
We present the results of solution-level scores in Figure 8. For solution-level scores, the tendencies discussed in §4.2 remain unchanged, but the distinction between the hypothesis and reference decreases, as the solution-level score can be disturbed by other steps. For reference-free evaluation, it adds difficulty in finding the specific step.
subfigure
subfigure
Appendix B Supplements for Evaluating LLMs
B.1 Location of the First Error Step
We examine the characteristics of the first error location identified by the validity score. We set the threshold to 0.25 for high recall. We present the main results in Figure 8. While there is significant noise in them, we can still observe general trends. For reasoning steps with incorrect results, the model is more prone to making mistakes at an early stage, leading to false results, especially when the lengths of reasoning steps are longer. For reasoning steps with correct results, the distribution is much flatter, indicating these errors hidden in reasoning steps can happen at any location.